Altira Corporation uses a periodic inventory system. The following information related to its merchandise inventory during the month of August 2018 is available:

Aug.1 Inventory on hand—3,500 units; cost $6.70 each.
8 Purchased 17,000 units for $6.20 each.
14 Sold 11,500 units for $12.70 each.
18 Purchased 9,500 units for $5.70 each.
25 Sold 10,500 units for $11.70 each.
31 Inventory on hand—8,000 units.


Required:
Determine the inventory balance Altira would report in its August 31, 2018, balance sheet and the cost of goods sold it would report in its August 2018 income statement using each of the following cost flow methods:

1. fifo

2. lifo

3. average cost
Answer #1
1-Aug inventory on hand 3500 6.7 23450
8-Aug purchased 17000 6.2 105400
18-Aug purchased 9,500 5.7 54150
total 30000 183000
Sold 22000
1) FIFO
cost of goods sold
1-Aug inventory on hand 3500 6.7 23450
8-Aug purchased 17000 6.2 105400
18-Aug purchased 1,500 5.7 8550
total 22000 137400
ending inventory 8000 *5.70= 45600
2) LIFO
18-Aug purchased 9,500 5.7 54150
8-Aug purchased 12,500 6.2 77500
total 22,000 131650
ending inventory
1-Aug inventory on hand 3500 6.7 23450
8-Aug purchased 4,500 6.2 27900
total 51350
3) Average cost
183000/30000
6.1
cost of goods sold   = 22000*6.10= 134200
Ending iventory      =   8,000*6.10= 48800
Answer-
COGS inventory
FIFO 137400 45600
LIFO 131,650 51350
Average cost 134200 48800